Solve Problems Using Linear Systems
16. A class trip is being planned. For one option, each student will pay $630. This includes two meals a day and accommodation for the 9-day trip. The other option offers three meals a day and accommodationfor the 9 days. This second option costs $720. What is the cost per meal? What is th cost per day for accommodation?
(The answer is $10 per meal, $50 per day for accommodation, but i dont know how to solve it?)

Answer by checkley77(12844)

9D+18M=630 MULTIPLY BY -1 & ADD THESE 2 EQUATIONS.
9D+27M=720
-9D-18M=-630
-----------------------
9M=90
M=90/9
M=$10 FOR EACH MEAL.
9D+18*10=630
9D+180=630
9D=630-180
9D=450
D=450/9
D=$50 PER DAY.